The length of a rectangular playing field is 5ft less than twice its width.If the perimeter of the playing field is 230ft,find the length and width of the field.
Answer by SandySharma(35)

Let the width of the field be x
Length = 2x -5
Perimeter = x + x + 2x - 5 + 2x - 5 = 6x - 10
Given that perimeter = 230 ft.
so, 4x - 10 = 230
Add 10 on both sides
6x = 230 + 10
6x = 240
Divide both sides by 6, you will get
x = 240/6 = 40
length = 2x - 5 = 2*40 - 5 = 80 - 5 = 75
Answer: length = 75ft. & width = 40ft.